A bullet of mass m = 40 g, moving horizontally with speed v, strikes a clay block of mass M=1.33 kg that is hanging on a light inextensible string of length L=0.767. The bullet becomes embedded in the block, which was originally at rest. What is the smallest value of vv which would cause the block-on-a-string to swing around and execute a complete vertical circle?
Answer #1

FONT _ Reference level Prom compervation mu = of momentum of system m +M) V E =A nee m +M Froom comervation of energy K»Eit PFONT _ Reference level Prom compervation mu = of momentum of system m +M) V E =A nee m +M Froom comervation of energy K»Eit PF.BD of bullet-clay system at top {m+m] 9 HT For circular motion (on tm) g +T = (mtm] vt for minimum speed, Teo .. (on+m) y =